OpenStudy (anonymous):

the volume of 10,000 drops of a liquid is 100 fluid ounces. What is the volume of 10 drops? I got 0.01

myininaya (myininaya):

I would do 100/1000=1/10=.1 how did you get .01

OpenStudy (anonymous):

divided 10 into 10,000

OpenStudy (anonymous):

no i mean 10 divided by 1,000

OpenStudy (anonymous):

No, divided 100 into 1000 = 0.1

OpenStudy (anonymous):

10,000 drops is 1,000 times of 10 drops

OpenStudy (anonymous):

So, you need to divide 100 ounces into 1000 = 0.1 ounce

OpenStudy (anonymous):

ok i see